OpenStudy (anonymous):

Find the exact value by using a half-angle identity. cosine of five pi divided by twelve.

OpenStudy (anonymous):

\[\frac{5\pi}{12}\]is half of \(\frac{5\pi}{6}\)

OpenStudy (anonymous):

so \[\cos(\frac{5pi}{12})=\sqrt{\frac{1+\cos(\frac{5\pi}{6})}{2}}\]

OpenStudy (anonymous):

that's my final answer?

OpenStudy (anonymous):

lol no

OpenStudy (anonymous):

your final answer should be a number find \(\cos(\frac{5\pi}{6})\) and plug it in

OpenStudy (anonymous):

\[\cos(\frac{5\pi}{6})=\frac{\sqrt3}{2}\] so you will have to do a bit of work to make that look reasonable

OpenStudy (anonymous):

oops my mistake \[\cos(\frac{5\pi}{6})=-\frac{\sqrt3}{2}\]

OpenStudy (anonymous):

\[\cos(\frac{5\pi}{12})=\sqrt{\frac{1-\frac{\sqrt3}{2}}{2}}\]

OpenStudy (xapproachesinfinity):

more simplified \[\cos(\frac{5\pi}{12})=\frac{\sqrt{2-\sqrt{3}}}{2}\]

OpenStudy (anonymous):

Thank you!!

OpenStudy (xapproachesinfinity):

welcome!

OpenStudy (xapproachesinfinity):

well thank @satellite73 though :)

OpenStudy (anonymous):

Soory, Thank you to each person for trying to help someone who knows nothing! You are both amazing!(:

OpenStudy (xapproachesinfinity):

haha welcome in place of everyone ^_^